ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 68 Rouillier, Fabrice, Loria, INRIA Lorrainne, France Real Solving Since Tarski, the existence of real solutions for systems of polynomial inequalities is known to be decidable. RealSolving is a long term project creating mathematical software which is devoted to the computation of real solutions to systems of polynomial equalities and inequalities using only exact computations. It has been used to solve problems in real algebraic geometry, algebra, computer vision, numerical methods, robotics and signal processing. This project received support from POSSO (Esprit BRA 6846) and FRISCO (LTR 21.024).